8. rs naturally in most ants. Soap solutions containing calcium hydroxide are used Formic acid occu to treat ant bites forming calcium formate. Given K of HC00H (formic acid) is 1.8 X 104 at room temperature, what answer is closest to the pH of the salt solution of 0.5 M calcium formate at room temperature? A) 2.02 B) 5.28 C) 8.72 D) 11.98 E) None of the them (A-D)
Answer #1

calcium formate is Ca(HCOO)2

so,

[HCOO-] = 2*[Ca(HCOO)2]

   = 2*0.5 M

   = 1.0 M

use:

Kb = Kw/Ka

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

Kb = (1.0*10^-14)/Ka

Kb = (1.0*10^-14)/1.8*10^-4

Kb = 5.556*10^-11

HCOO- dissociates as

HCOO- + H2O -----> HCOOH + OH-

0.5 0 0

0.5-x x x

Kb = [HCOOH][OH-]/[HCOO-]

Kb = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Kb = x*x/(c)

so, x = sqrt (Kb*c)

x = sqrt ((5.556*10^-11)*0.5) = 5.27*10^-6

since c is much greater than x, our assumption is correct

so, x = 5.27*10^-6 M

use:

pOH = -log [OH-]

= -log (5.27*10^-6)

= 5.2782

use:

PH = 14 - pOH

= 14 - 5.2782

= 8.7218

Answer: 8.72

Answer: C